# Human as Points: Explicit Point-based 3D Human Reconstruction from Single-view RGB Images

### v2

@@ III-B2 SMPL Estimation and Rectification
Let $\mathcal{S}_{v}=(\mathcal{V}_{v},~{}\mathcal{F}_{v})$ and $\overline{\mathcal{S}}_{v}=(\overline{\mathcal{V}}_{v},~{}\overline{\mathcal{F}}_{v})$ be the visible and the invisible parts of $\mathcal{S}$ (i.e. $\mathcal{S}=\mathcal{S}_{v}\cup\overline{\mathcal{S}}_{v}$), respectively, where $\mathcal{V}_{v}$, $\overline{\mathcal{V}}_{v}$, $\mathcal{F}_{v}$, and ¯$\overline{\mathcal{F}}_{v}$ stand for the vertex and face sets of the two parts, and $\widehat{\mathcal{S}}$ the rectified SMPL model. Generally, this problem can be achieved through iterative rectifications to the shape parameters $\bm{\beta}$ and pose parameters $\bm{\theta}$ of $\mathcal{S}$ under the objective of minimizing the distance between $\mathcal{S}_{v}$ and $\mathcal{P}$. Besides, to avoid any unwanted scenarios where the invisible part gets unnaturally close to $\mathcal{P}$, resulting in unnatural poses, we add a regularization term to maximize the distance between $\mathcal{P}$ and $\overline{\mathcal{S}}_{v}$. We also regularize shape parameters $\bm{\beta}$ due to their inherent zero-mean nature as established [5]. Moreover, the optimization is subject to two constraints: (1) the difference between the 2D keypoints of $\widehat{\mathcal{S}}$ and $\mathcal{S}$, preventing excessive variations in the pose of $\widehat{\mathcal{S}}$ on the 2D image, and (2) the silhouette difference between the mask of $\widehat{\mathcal{S}}$ (i.e., $\mathcal{M}_{\widehat{\mathcal{S}}}$) when projected onto the 2D image and the human mask from $\mathcal{I}$ (i.e., $\mathcal{M}_{h}$), preserving the overall human shape of $\widehat{\mathcal{S}}$. In all, we explicitly formulate the problem as
$\operatorname*{{\color[rgb]{0,0,0}min}}_{\bm{{\color[rgb]{0,0,0}\beta}},\bm{{\color[rgb]{0,0,0}\theta}}}~{}{\color[rgb]{0,0,0}\mathcal{L}_{\mathrm{r1}}}{\color[rgb]{0,0,1}=}$ ${\color[rgb]{0,0,0}\lambda_{1}\texttt{P2F}(\mathcal{P},~{}\mathcal{S}_{v})+\lambda_{2}\texttt{CD}(\mathcal{V}_{v},\mathcal{P})}-$ (1)
@@ IV-A Data Preparation
One of the primary development bottlenecks in the research area of single-view human reconstruction is the issue of data insufficiency. To address this issue, we have created a new dataset, referred to as CityUHuman [^2] https://drive.google.com/drive/folders/1N-0UeC4_oU4R5lrn5avn_oZvaNp
m1uBG?usp=sharing
, which expands the number of available human body scans. We will make this dataset publicly available for research purposes so that it serves as a valuable supplement for assessing and benchmarking the performance of single-view human reconstruction methods. In comparison to the Thuman dataset [54, 66], the human models in the CityUHuman dataset exhibit smoother and more accurate surface details. As shown in Fig. 7, the Thuman2 scan displays a disrupted geometry at the pocket area due to its rough surface, whereas the CityUHuman scan preserves the finer geometric details more accurately.
@@ IV-C Comparisons with State-of-the-Art Methods
Finally, to assess the effectiveness of our HaP in real-world scenarios, we gathered a collection of images featuring diverse clothing styles and poses from the internet to infer their full 3D structures. Fig. 10 visually compares the results of our HaP with ICON and ECON, where it can be seen that HaP can correctly reconstruct the clothing details and poses, while ICON cannot reconstruct the loose clothing details and ECON fails to recover correct human poses. More results of our HaP are depicted in Fig. 11, demonstrating commendable performance in realistic scenarios. We also refer readers to the video demo [^5] https://youtu.be/x8OtDLgkBHU.
TABLE III: Comparison of different SMPL rectification methods. Thuman2 is used for evaluation.
@@
IV-D Analysis of SMPL Rectification
@@
IV-D1 Defects of rendering-based optimization
@@
In addition to optimizing SMPL models with normal maps, we further experimented with depth maps. The qualitative and quantitative comparisons are presented in Fig. 13 and Table III, respectively.
@@
IV-D2 The effectiveness of optimization in 3D space
Several methods [60, 4] have evaluated the efficacy of registering SMPL models within the implicit 3D space. In the context of monocular human reconstruction, PaMIR [68] also optimizes the SMPL models in the implicit 3D space by minimizing the occupancy loss to make the SMPL align with the predicted implicit function, as shown in Table III and Fig. 13, it achieves better quantitative and qualitative results than rendering-based methods.
@@ IV-D2 The effectiveness of optimization in 3D space
TABLE IV: The ablation studies on different modules of HaP. “SR” and “DR” are the SMPL rectification and depth replacement modules, respectively. Thuman3.0 is used for evaluation. The row highlighted in blue corresponds to our final HaP. $\downarrow$” means the lower, the better.
@@ IV-E Ablation Studies
IV-E1 Depth estimation
Note that MIM [62] achieves more accurate depth on Thuman3.0 than P3Depth [43], i.e., 0.0158 vs. 0.0189 in terms of RMSE. From the $1^{st}$ row of Table IV, as expected, a better depth estimation method can lead to better reconstruction performance.
@@ IV-E1 Depth estimation
IV-E2 Conditioned diffusion-based generation
We replaced PNet [34] with VRCNet [41], a point cloud completion network, to evaluate the effectiveness of the diffusion-based generation network.
As illustrated in Fig. 15, VRCNet generates highly noisy point clouds, which adversely affects the accuracy of the depth replacement module and leads to poor performance of HaP, as indicated in the $2^{nd}$ row of Table IV, with a higher normal error. Moreover, our diffusion-based point cloud generation is able to effectively remove the noisy points produced by depth estimation (see the noisy points between legs in Fig. 16).
IV-E3 SMPL estimation and rectification
Compared to PyMAF [67], PIXIE [16] (the $3^{rd}$ row) slightly under-performs in terms of CD and Normal metrics. As listed in the 4th row of Table IV, although removing the SMPL rectification module does not drastically reduce the numerical results, the head looks rather weird and unnatural, as exemplified in Fig. 17.
Moreover, we conducted experiments by replacing the SMPL models used in ICON and ECON with our HaP-rectified SMPL models. The modified methods are denoted as ICON-S and ECON-S. As shown in Table V, a comparison between ICON with ICON-S and ECON with ECON-S reveals that HaP-refined SMPL models can improve the performance of ICON and ECON, demonstrating the effectiveness of SMPL rectification. Besides, Table V shows that HaP consistently outperforms ICON-S and ECON-S, underscoring the effectiveness of our explicit diffusion-based framework over the implicit function models utilized by ICON and ECON.
IV-E4 Refinement stage
As listed in the $5^{th}$ row of Table IV, although HaP without the depth replacement strategy achieves satisfactory performance in terms of CD and P2F, there is a significant increase in the Normal error, demonstrating its lack of effectiveness in recovering surface details.
Furthermore, we assess the impact of omitting the displacement-based refinement module, $\texttt{PNet}{\Theta_{2}}$, which improves the point cloud’s geometric details by learning point-wise displacements. As shown in the $6^{th}$ and $7^{th}$ rows of Table IV, the substantial performance drop across all metrics when $\texttt{PNet}{\Theta_{2}}$ is removed highlights its essential role in producing high-quality results.
TABLE V: Quantitative comparisons of ICON, ICON-S, ECON, ECON-S, and our Hap on Thuman3 when the same SMPL pose parameters obtained by our HaP were adopted by all methods.
@@ IV-F Perceptual Study
In the first perceptual evaluation, the questionnaire [^6] Perceptual Evaluation 1: https://forms.gle/23Hv4cDGbki56w7Z7 consists of ten questions. Each question presented an RGB image and six videos. The first video in each question displayed the ground truth scanned model, while the remaining five videos corresponded to five different methods. The order of the methods was randomized across the ten questions. Participants were asked to rate the methods on a scale from 1 to 5 based on their similarity to the real model. For instance, if method A was judged as the most accurate, it would receive a score of 5; if it was considered the least accurate, it would receive a score of 1. An intuitive guide for scoring was provided [^7] Intuitive Figure: https://forms.gle/M9X2EiyrJ7EfXp6d8. Participants were not filtered via catch trials, and no special training was provided. The evaluation focused on aspects such as posture accuracy, limb completeness, clothing style reconstruction, rationality of the non-visible side, and surface smoothness. As shown in Fig. 18 (a), HaP achieved the highest mean score.
In the second perceptual evaluation [^8] Perceptual Evaluation 2: https://forms.gle/kzybD19NronWAiec6, participants were shown the same videos and asked to select the best method between IntegratedPIFu and HaP, as they had obtained the highest scores in the first evaluation. Fig. 18 (b) shows that HaP was selected as the best method in six out of the ten samples.
IV-G Discussion on Limitations

### v3

@@ II-D Clothed Human Reconstruction
II-E 3D Generative Models
Significant progress has been made in advancing deep generative models for 3D content creation. The representative optimization-based approaches such as DreamFusion [49] and Magic123 [51], leverage pre-trained 2D image diffusion models by score distillation, which often struggle with geometric inconsistency across views. DreamCraft3D [61] introduces a hierarchical 3D generation framework that ensures geometric consistency and high-fidelity textures, incorporating bootstrapped score distillation to enhance appearance quality. Huang et al. [23] proposed a text-to-3D model utilizing normal-adapted and depth-adapted diffusion models for detailed geometry generation, alongside a normal-aligned diffusion model to improve appearance realism. Li et al. developed PSHuman [30], which employs a diffusion model to generate multi-view images and normal maps conditioned on an SMPL-X prior, followed by explicit human body surface reconstruction.
Some methods aim to generate 3D contents directly in 3D space. Ren et al. [55] introduced a hierarchical voxel-diffusion generative model capable of producing large-scale 3D objects. Hunyuan3D [75] learns a latent representation of 3D objects and employs DiT [48] to train a generative model that supports 3D painting. Unlike the above-mentioned 3D generative methods, HaP learns a conditional diffusion model and manipulates point clouds entirely in 3D space rather than 2D image space, providing a more intuitive and direct learning and manipulation process for human body reconstruction.
@@ III-B2 SMPL Estimation and Rectification
Let $\mathcal{S}_{v}=(\mathcal{V}_{v},~{}\mathcal{F}_{v})$ and $\overline{\mathcal{S}}_{v}=(\overline{\mathcal{V}}_{v},~{}\overline{\mathcal{F}}_{v})$ be the visible and the invisible parts of $\mathcal{S}$ (i.e. $\mathcal{S}=\mathcal{S}_{v}\cup\overline{\mathcal{S}}_{v}$), respectively, where $\mathcal{V}_{v}$, $\overline{\mathcal{V}}_{v}$, $\mathcal{F}_{v}$, and ¯$\overline{\mathcal{F}}_{v}$ stand for the vertex and face sets of the two parts, and $\widehat{\mathcal{S}}$ the rectified SMPL model. Generally, this problem can be achieved through iterative rectifications to the shape parameters $\bm{\beta}$ and pose parameters $\bm{\theta}$ of $\mathcal{S}$ under the objective of minimizing the distance between $\mathcal{S}_{v}$ and $\mathcal{P}$.
Besides, to avoid any unwanted scenarios where the invisible part gets unnaturally close to $\mathcal{P}$, resulting in unnatural poses, we add a regularization term to maximize the distance between $\mathcal{P}$ and $\overline{\mathcal{S}}_{v}$. We also regularize shape parameters $\bm{\beta}$ due to their inherent zero-mean nature as established [5]. Moreover, the optimization is subject to two constraints: (1) the difference between the 2D keypoints of $\widehat{\mathcal{S}}$ and $\mathcal{S}$, preventing excessive variations in the pose of $\widehat{\mathcal{S}}$ on the 2D image, and (2) the silhouette difference between the mask of $\widehat{\mathcal{S}}$ (i.e., $\mathcal{M}_{\widehat{\mathcal{S}}}$) when projected onto the 2D image and the human mask from $\mathcal{I}$ (i.e., $\mathcal{M}_{h}$), preserving the overall human shape of $\widehat{\mathcal{S}}$. In all, we explicitly formulate the problem as
$\operatorname*{{\color[rgb]{0,0,0}min}}_{\bm{{\color[rgb]{0,0,0}\beta}},\bm{{\color[rgb]{0,0,0}\theta}}}~{}{\color[rgb]{0,0,0}\mathcal{L}_{\mathrm{r1}}}{\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}\lambda_{1}\texttt{P2F}(\mathcal{P},~{}\mathcal{S}_{v})+\lambda_{2}\texttt{CD}(\mathcal{V}_{v},\mathcal{P})}-$ (1)
@@ IV-A Data Preparation
One of the primary development bottlenecks in the research area of single-view human reconstruction is the issue of data insufficiency. To address this issue, we have created a new dataset, referred to as CityUHuman [^2] https://github.com/yztang4/HaP, which expands the number of available human body scans. We will make this dataset publicly available for research purposes so that it serves as a valuable supplement for assessing and benchmarking the performance of single-view human reconstruction methods. In comparison to the Thuman dataset [60, 73], the human models in the CityUHuman dataset exhibit smoother and more accurate surface details. As shown in Fig. 7, the Thuman2 scan displays a disrupted geometry at the pocket area due to its rough surface, whereas the CityUHuman scan preserves the finer geometric details more accurately.
@@ IV-C Comparisons with State-of-the-Art Methods
Finally, to assess the effectiveness of our HaP in real-world scenarios, we gathered a collection of images featuring diverse clothing styles and poses from the internet to infer their full 3D structures. Fig. 10 visually compares the results of our HaP with ICON and ECON, where it can be seen that HaP can correctly reconstruct the clothing details and poses, while ICON cannot reconstruct the loose clothing details and ECON fails to recover correct human poses. More results of our HaP are depicted in Fig. 11, demonstrating commendable performance in realistic scenarios. We also refer readers to the hapresultsvideodemo.mp4 in the supplementary material.
Figure 13: Comparison with generative models. (a) Hunyuan3D 2.0 [75]. (b) PSHuman [30]. (c) Ours.
TABLE III: Quantitative comparisons of Hunyuan3D 2.0 [75], PSHuman [30] and our Hap on Thuman3.
CD P2F Normal
Hunyuan3D 1.267 1.229 2.985
PSHuman 1.045 1.171 2.386
HaP 0.860 0.925 2.170
IV-D Comparisons with 3D Generative Methods
We also compared HaP with two 3D generative methods: Hunyuan3D 2.0 [75], a general-purpose approach, and PSHuman [30], a human-centric method. Note that both models were trained with large-scale datasets, whereas HaP, specifically for generating 3D human bodies, was trained only with 500 scans from Thuman2.0 [73]. As shown in Table III, HaP achieves the best overall quantitative performance.
In addition to the numerical results, we conducted a visual comparison to further analyze the strengths and limitations of each method. As illustrated in Fig. 13, PSHuman [30] occasionally incorrectly generates closed-leg postures instead of the expected open-leg stance, highlighting its limitations in pose estimation. In contrast, HaP accurately preserves such structural details. Furthermore, Hunyuan3D 2.0 [75] sometimes struggles to generate complete body parts, particularly under severe occlusions. This limitation suggests that relying solely on data-driven learning without incorporating strong human priors can hinder the generation of complete human bodies.
TABLE IV: Comparison of different SMPL rectification methods. Thuman2 is used for evaluation.
@@ IV-D Comparisons with 3D Generative Methods
IV-E Analysis of SMPL Rectification
@@
IV-E1 Defects of rendering-based optimization
@@
In addition to optimizing SMPL models with normal maps, we further experimented with depth maps. The qualitative and quantitative comparisons are presented in Fig. 14 and Table IV, respectively.
@@
IV-E2 The effectiveness of optimization in 3D space
Several methods [67, 4] have evaluated the efficacy of registering SMPL models within the implicit 3D space. In the context of monocular human reconstruction, PaMIR [76] also optimizes the SMPL models in the implicit 3D space by minimizing the occupancy loss to make the SMPL align with the predicted implicit function, as shown in Table IV and Fig. 14, it achieves better quantitative and qualitative results than rendering-based methods.
@@ IV-E2 The effectiveness of optimization in 3D space
TABLE V: The ablation studies on different modules of HaP. “SR” and “DR” are the SMPL rectification and depth replacement modules, respectively. Thuman3.0 is used for evaluation. The row highlighted in blue corresponds to our final HaP. $\downarrow$” means the lower, the better.
@@ IV-F Ablation Studies
IV-F1 Depth estimation
Note that MIM [69] achieves more accurate depth on Thuman3.0 than P3Depth [45], i.e., 0.0158 vs. 0.0189 in terms of RMSE. From the $1^{st}$ row of Table V, as expected, a better depth estimation method can lead to better reconstruction performance.
@@ IV-F1 Depth estimation
IV-F2 Conditioned diffusion-based generation
We replaced PNet [36] with VRCNet [43], a point cloud completion network, to evaluate the effectiveness of the diffusion-based generation network. As illustrated in Fig. 16, VRCNet generates highly noisy point clouds, which adversely affects the accuracy of the depth replacement module and leads to poor performance of HaP, as indicated in the $2^{nd}$ row of Table V, with a higher normal error. Moreover, our diffusion-based point cloud generation is able to effectively remove the noisy points produced by depth estimation (see the noisy points between legs in Fig. 17).
IV-F3 SMPL estimation and rectification
Compared to PyMAF [74], PIXIE [16] (the $3^{rd}$ row) slightly under-performs in terms of CD and Normal metrics. As listed in the 4th row of Table V, although removing the SMPL rectification module does not drastically reduce the numerical results, the head looks rather weird and unnatural, as exemplified in Fig. 18.
Moreover, we conducted experiments by replacing the SMPL models used in ICON and ECON with our HaP-rectified SMPL models. The modified methods are denoted as ICON-S and ECON-S. As shown in Table VI, a comparison between ICON with ICON-S and ECON with ECON-S reveals that HaP-refined SMPL models can improve the performance of ICON and ECON, demonstrating the effectiveness of SMPL rectification. Besides, Table VI shows that HaP consistently outperforms ICON-S and ECON-S, underscoring the effectiveness of our explicit diffusion-based framework over the implicit function models utilized by ICON and ECON.
IV-F4 Refinement stage
As listed in the $5^{th}$ row of Table V, although HaP without the depth replacement strategy achieves satisfactory performance in terms of CD and P2F, there is a significant increase in the Normal error, demonstrating its lack of effectiveness in recovering surface details.
Furthermore, we assess the impact of omitting the displacement-based refinement module, $\texttt{PNet}{\Theta_{2}}$, which improves the point cloud’s geometric details by learning point-wise displacements. As shown in the $6^{th}$ and $7^{th}$ rows of Table V, the substantial performance drop across all metrics when $\texttt{PNet}{\Theta_{2}}$ is removed highlights its essential role in producing high-quality results.
TABLE VI: Quantitative comparisons of ICON, ICON-S, ECON, ECON-S, and our Hap on Thuman3 when the same SMPL pose parameters obtained by our HaP were adopted by all methods.
@@ IV-G Perceptual Study
In the first perceptual evaluation, the questionnaire (Perceptual_Evaluation_1.pdf in the supplementary material) consists of ten questions. Each question presented an RGB image and six videos. The first video in each question displayed the ground truth scanned model, while the remaining five videos corresponded to five different methods. The order of the methods was randomized across the ten questions. Participants were asked to rate the methods on a scale from 1 to 5 based on their similarity to the real model. For instance, if method A was judged as the most accurate, it would receive a score of 5; if it was considered the least accurate, it would receive a score of 1. An intuitive guide for scoring was provided (Intuitive_Figure.pdf in the supplementary material). Participants were not filtered via catch trials, and no special training was provided. The evaluation focused on aspects such as posture accuracy, limb completeness, clothing style reconstruction, rationality of the non-visible side, and surface smoothness. As shown in Fig. 19 (a), HaP achieved the highest mean score.
In the second perceptual evaluation (Perceptual_Evaluation_2.pdf in the supplementary material), participants were shown the same videos and asked to select the best method between IntegratedPIFu and HaP, as they had obtained the highest scores in the first evaluation. Fig. 19 (b) shows that HaP was selected as the best method in six out of the ten samples.
IV-H Discussion on Limitations